# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v1

@@ A.5 Understanding PreDeToR
A.6 Empirical Study: Increasing number of Actions
@@
(1) The PreDeToR (-$\tau$) exploits the underlying latent structure and reward correlation between the different tasks even when $A>n$ but $M_{\text{pre}}\gg A$.
(2) The PreDeToR (-$\tau$) has lower regret than its weak demonstrator even when it is trained on $\mathcal{H}_{\text{train}}\subseteq\mathcal{D}_{\text{pre}}$.
(3) The PreDeToR (-$\tau$) performance drops for complex non-linear bandit setting with increasing number of dimensions.
@@
Experimental Result: We observe these outcomes in Figure 6. In Figure 6(a) we show the linear bandit setting for $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=100$, $n=50$ and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy and AD. Note that for any task $m$ the Thomp will not be able to sample all the actions even once. The weak performance of DPT-greedy can be attributed to both short horizons and the inability to estimate the optimal action for such a short horizon $n<A$. The AD performs similar to the demonstrator Thomp because of its training. Observe that PreDeToR (-$\tau$) has similar regret to LinUCB and lower regret than Thomp which also shows that PreDeToR is exploiting the latent linear structure of the underlying tasks. In Figure 6(b) we show the non-linear bandit setting for horizon $n=40$, $M_{\text{pre}}=200000$, $A=60$, $d=2$, and $|\mathcal{A}^{\text{inv}}|=5$. The demonstrator $\pi^{w}$ is the Thomp algorithm. The training of AD ensures that it does not outperform the demonstrator Thomp. Again we observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD and LinUCB which fails to perform well in this non-linear setting due to its algorithmic design.
A.7 Empirical Study: Increasing Horizon
In this section, we discuss the performance of PreDeToR with respect to an increasing horizon for each task $m\in[M]$. However, note that the number of tasks $M_{\text{pre}}\geq n$. Note that Lee et al. [2023] studied linear bandit setting for $n=200$. We study the setting up to a similar horizon scale.
@@ A.7 Empirical Study: Increasing Horizon
(1) The PreDeToR (-$\tau$) learns to exploit the underlying latent structure and reward correlation between the different tasks even when $A\geq n$, $M_{\text{pre}}\gg A$ with increasing horizon.
(2) The PreDeToR (-$\tau$) outperforms its weak demonstrator even when trained on $\mathcal{H}_{\text{train}}\subseteq\mathcal{D}_{\text{pre}}$.
@@
Experimental Result: We observe these outcomes in Figure 7. In Figure 7 we show the linear bandit setting for $M_{\text{pre}}=150000$, $M_{\text{test}}=200$, $A=20$, $n=\{20,40,60,100,120,140,200\}$ and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Observe from Figure 7(a), 7(b), 7(c), Figure 7(d), 7(e), 7(f) and 7(g) that PreDeToR (-$\tau$) is closer to LinUCB and outperforms Thomp which also shows that PreDeToR (-$\tau$) is learning the latent linear structure of the underlying tasks. In Figure 7(h) we plot the regret of all the baselines with respect to the increasing horizon. Again we see that PreDeToR (-$\tau$) is closer to LinUCB and outperforms DPT-greedy, AD and Thomp. The training of AD ensures that it does not outperform the demonstrator Thomp. This shows that PreDeToR (-$\tau$) is able to exploit the latent structure and reward correlation across the tasks for varying horizon length.
A.8 Empirical Study: Increasing Dimension
@@
(1) The PreDeToR (-$\tau$) learns to exploit the underlying latent structure and reward correlation between the different tasks even when $A\geq n$ but $M_{\text{pre}}\gg A$.
(2) The PreDeToR (-$\tau$) has lower regret than its weak demonstrator even when trained on $\mathcal{D}_{\text{train}}\subseteq\mathcal{D}_{\text{pre}}$.
(3) The PreDeToR (-$\tau$) has lower regret than LinUCB for larger dimension.
@@
Experimental Result: We observe these outcomes in Figure 7. In Figure 7 we show the linear bandit setting for horizon $n=20$, $M_{\text{pre}}=160000$, $M_{\text{test}}=200$, $A=20$, and $d=\{10,20,30,40\}$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Observe from Figure 8(a), 8(b), 8(c), and 8(d) that PreDeToR (-$\tau$) is closer to LinUCB and has lower regret than Thomp which also shows that PreDeToR (-$\tau$) is exploiting the latent linear structure of the underlying tasks. In Figure 8(e) we plot the regret of all the baselines with respect to the increasing dimension. Again we see that PreDeToR (-$\tau$) has lower regret than DPT-greedy, AD and Thomp. Observe that with increasing dimension PreDeToR is able to outperform LinUCB. This shows that the PreDeToR (-$\tau$) is able to exploit reward correlation across tasks for varying dimensions.
A.9 Empirical Study: Increasing Attention Heads
@@
(1) The PreDeToR (-$\tau$) learns to exploit the underlying latent structure and reward correlation between the different tasks even when $A\geq n$ but $M_{\text{pre}}\gg A$.
(2) The PreDeToR (-$\tau$) has lower regret than its weak demonstrator even when trained on $\mathcal{D}_{\text{train}}\subseteq\mathcal{D}_{\text{pre}}$.
(3) The PreDeToR (-$\tau$) has lower regret compared to DPT-greedy, AD with increasing attention heads.
@@
Experimental Result: We observe these outcomes in Figure 9. In Figure 9 we show the non-linear bandit setting for horizon $n=20$, $M_{\text{pre}}=160000$, $M_{\text{test}}=200$, $A=20$, $\mathrm{heads}=\{2,4,6,8\}$ and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions atmost once. Observe from Figure 9(a), 9(b), 9(c), and 9(d) that PreDeToR (-$\tau$) has lower regret than AD, Thomp and LinUCB which also shows that PreDeToR (-$\tau$) is exploiting the latent linear structure of the underlying tasks for the non-linear setting. The training of AD ensures that it does not outperform the demonstrator Thomp. In Figure 9(f) we plot the regret of all the baselines with respect to the increasing attention heads. Again we see that PreDeToR (-$\tau$) regret decreases as we increase the attention heads.
A.10 Empirical Study: Increasing Number of Tasks
@@
(1) The PreDeToR (-$\tau$) fails to exploit the underlying latent structure and reward correlation when the number of tasks is small.
(2) As the number of tasks increases the PreDeToR (-$\tau$) learns the latent structure and the underlying reward correlation.
(3) The PreDeToR (-$\tau$) has lower regret compared to DPT-greedy, AD with increasing number of tasks.
@@
Experimental Result: We observe these outcomes in Figure 10. In Figure 10 we show the linear bandit setting for horizon $n=20$, $M_{\text{pre}}\in\{5000,10000,50000,100000,150000\}$, $M_{\text{test}}=200$, $A=20$, and $d=40$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$), AD and DPT-greedy suffer more regret than the LinUCB when the number of tasks is small ($M_{\text{train}}\in\{5000,10000\}$ in Figure 10(a), and 10(b). However in Figure 10(c), 10(d), 10(e), and 10(f) we show that PreDeToR has lower regret than Thomp and matches LinUCB. This shows that PreDeToR (-$\tau$) is exploiting the latent linear structure of the underlying tasks for the non-linear setting. Moreover, observe that as $M_{\text{train}}$ increases the PreDeToR has lower cumulative regret than DPT-greedy, AD. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Therefore DPT-greedy does not perform as well as PreDeToR. The training of AD ensures that it does not outperform the demonstrator Thomp. Finally, note that the result shows that PreDeToR (-$\tau$) is able to exploit the reward correlation across the tasks better as the number of tasks increases.
A.11 Exploration of PreDeToR(-$\tau$)
In this section, we discuss the exploration of PreDeToR in the linear bandit setting discussed in Section 4. Recall that the linear bandit setting consist of horizon $n=25$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=10$, and $d=2$. The demonstrator $\pi^{w}$ is the Thomp algorithm and we observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD and matches the performance of LinUCB. Therefore PreDeToR (-$\tau$) behaves almost optimally in this setting and so we analyze how PreDeToR conducts exploration for this setting.
(1) The PreDeToR (-$\tau$) has a two phase exploration. In the first phase it explores with a strong prior over the training data. In the second phase, once the task data has been observed for a few rounds it switches to task-based exploration.
(2) PreDeToR (-$\tau$) samples the optimal action in each task more than DPT-greedy. PreDeToR(-$\tau$) considers a larger set of possible optimal action candidate set than DPT-greedy.
Then in Figure 11(b) we show how the sampling distribution of DPT-greedy, PreDeToR  and PreDeToR-$\tau$ change in the first $10$ and last $10$ rounds for all the tasks where action $5$ is optimal. To plot this graph we first sum over the individual pulls of the action taken by each algorithm over the first $10$ and last $10$ rounds. Then we average these counts over all test tasks where action $5$ is optimal. From the figure Figure 11(b) we see that PreDeToR(-$\tau$) consistently pulls the action $5$ more than DPT-greedy. It also explores other optimal actions like $\{2,3,6,7,10\}$ but discards them quickly in favor of the optimal action $5$ in these tasks. This shows that PreDeToR (-$\tau$) only considers the optimal actions seen from the training data. Once sufficient observation have been observed for the task it switches to task-based exploration and samples the optimal action more than DPT-greedy.
Finally, we plot the feasible action set considered by DPT-greedy, PreDeToR, and PreDeToR-$\tau$ in Figure 11(c). To plot this graph again we consider the test tasks where the optimal action is $5$. Then we count the number of distinct actions that are taken from round $t$ up until horizon $n$. Finally we average this over all the considered tasks where the optimal action is $5$. We call this the candidate action set considered by the algorithm. From the Figure 11(c) we see that DPT-greedy explores the least and gets stuck with few actions quickly (by round $10$). Note that the actions DPT-greedy samples are sub-optimal and so it suffers a high cumulative regret (see Figure 1(a)). PreDeToR explore slightly more than DPT-greedy, but PreDeToR-$\tau$ explores the most.
(a) Train Optimal Action Distribution
(b) Distribution of action sampling in all tasks where action $5$ is optimal
Figure 11: Exploration Analysis of PreDeToR(-$\tau$)
@@ A.12 Exploration of PreDeToR(-$\tau$) in New Arms Setting
(1) The PreDeToR (-$\tau$) is robust to changes when the number of changing arms is small. It samples the optimal action in each task more than DPT-greedy.
(2) When the number of in-variant actions is small then all algorithms PreDeToR, PreDeToR-$\tau$, and AD sample the optimal action very less. This shows that the latent shared structure across the tasks is very important.
@@
Then in Figure 12(b) we show how the sampling distribution of DPT-greedy, PreDeToR  and PreDeToR-$\tau$ change in the first $10$ and last $10$ rounds for all the tasks where action $17$ is optimal. We plot this graph the same way as discussed in Section A.11. From the figure Figure 12(b) we see that PreDeToR(-$\tau$) consistently pulls the action $17$ more than DPT-greedy. It also explores other optimal actions like $\{1,2,3,8,9,15\}$ but discards them quickly in favor of the optimal action $17$ in these tasks.
@@
We now consider the setting where the number of invariant actions is $|\mathcal{A}^{\text{inv}}|=15$. We again show in Figure 14(a) the training distribution of the optimal actions. For each bar, the frequency indicates the number of tasks where the action (shown in the x-axis) is the optimal action. Then in Figure 14(b) we show how the sampling distribution of DPT-greedy, PreDeToR  and PreDeToR-$\tau$ change in the first $10$ and last $10$ rounds for all the tasks where action $17$ is optimal. We plot this graph the same way as discussed in Section A.11. From the figure Figure 14(b) we see that none of the algorithms PreDeToR, PreDeToR-$\tau$, DPT-greedy consistently pulls the action $17$ more than other actions. This shows that the common underlying actions across the tasks matter for learning the epxloration.
@@
A.13 Data Collection Analysis
In this section, we analyze the performance of PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LinUCBwhen the weak demonstrator $\pi^{w}$ is Thomp, LinUCB, or Uniform. We again consider the linear bandit setting discussed in Section 4. Recall that the linear bandit setting consist of horizon $n=25$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=10$, and $d=2$. Finally, we show the cumulative regret by the above baselines in Figure 16(a), 16(b), and 16(b) when data is collected through Thomp, LinUCB, and Uniform respectively.
Outcomes: We first discuss the main outcomes of our experimental results for different data collection:
(1) The PreDeToR (-$\tau$) excels in exploiting the underlying latent structure and reward correlation when the data diversity is high.
(2) The PreDeToR (-$\tau$) fails to exploit the underlying latent structure and reward correlation when the data diversity is less.
(3) The PreDeToR (-$\tau$) has lower regret compared to DPT-greedy in all settings. AD outperforms PreDeToR (-$\tau$) when data is collected by LinUCB.
(a) Thomp data collection
(b) LinUCB data collection
(c) Uniform data collection
A.14 Empirical Validation of Theoretical Result
@@
(1) The PreDeToR (-$\tau$) excels in predicting the rewards for test tasks when the number of training (source) tasks is large.
(2) The PreDeToR (-$\tau$) has a poor prediction of the rewards for test tasks when the number of training (source) tasks is small.
(a) Prediction Error for $100000$ tasks
(b) Prediction Error for $200000$ tasks
@@
A.15 Empirical Study: Offline Performance
@@
(1) The PreDeToR performs comparably with respect to cumulative regret against DPT-greedy both for the long horizon and the short horizon setting.
(2) The PreDeToR performs comparably with respect to cumulative regret against DPT-greedy both for the linear and the non-linear bandit setting.
(3) The PreDeToR and DPT-greedy has comparable regret to the weak demonstrator Thomp and LinUCB in the offline setting.

### v2

@@ A.6 Connection between PreDeToR and Linear Multivariate Gaussian Model
A.7 Empirical Study: Increasing number of Actions
@@
Finding 11: PreDeToR (-$\tau$) outperforms DPT-greedy, and AD, even when $A>n$ but $M_{\text{pre}}\gg A$.
@@
Experimental Result: We observe these outcomes in Figure 11. In Figure 11(a) we show the linear bandit setting for $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=100$, $n=50$ and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy and AD. Note that for any task $m$ the Thomp will not be able to sample all the actions even once. The weak performance of DPT-greedy can be attributed to both short horizons and the inability to estimate the optimal action for such a short horizon $n<A$. The AD performs similar to the demonstrator Thomp because of its training. Observe that PreDeToR (-$\tau$) has similar regret to LinUCB and lower regret than Thomp which also shows that PreDeToR is exploiting the latent linear structure of the underlying tasks. In Figure 11(b) we show the non-linear bandit setting for horizon $n=40$, $M_{\text{pre}}=200000$, $A=60$, $d=2$, and $|\mathcal{A}^{\text{inv}}|=5$. The demonstrator $\pi^{w}$ is the Thomp algorithm. Again we observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD and LinUCB which fails to perform well in this non-linear setting due to its algorithmic design.
A.8 Empirical Study: Increasing Horizon
In this section, we discuss the performance of PreDeToR with respect to an increasing horizon for each task $m\in[M]$. However, note that the number of tasks $M_{\text{pre}}\geq n$. Note that Lee et al. (2023) studied linear bandit setting for $n=200$. We study the setting up to a similar horizon scale.
@@ A.8 Empirical Study: Increasing Horizon
Finding 12: PreDeToR (-$\tau$) outperforms DPT-greedy, and AD with increasing horizon.
@@
Experimental Result: We observe these outcomes in Figure 12. In Figure 12 we show the linear bandit setting for $M_{\text{pre}}=150000$, $M_{\text{test}}=200$, $A=20$, $n=\{20,40,60,100,120,140,200\}$ and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Observe from Figure 12(a), 12(b), 12(c), Figure 12(d), 12(e), 12(f) and 12(g) that PreDeToR (-$\tau$) is closer to LinUCB and outperforms Thomp which also shows that PreDeToR (-$\tau$) is learning the latent linear structure of the underlying tasks. In Figure 12(h) we plot the regret of all the baselines with respect to the increasing horizon. Again we see that PreDeToR (-$\tau$) is closer to LinUCB and outperforms DPT-greedy, AD and Thomp. This shows that PreDeToR (-$\tau$) is able to exploit the latent structure and reward correlation across the tasks for varying horizon length.
A.9 Empirical Study: Increasing Dimension
@@
Finding 13: PreDeToR (-$\tau$) outperforms DPT-greedy, AD with increasing dimension and has lower regret than LinUCB for larger dimension.
@@
Experimental Result: We observe these outcomes in Figure 12. In Figure 12 we show the linear bandit setting for horizon $n=20$, $M_{\text{pre}}=160000$, $M_{\text{test}}=200$, $A=20$, and $d=\{10,20,30,40\}$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Observe from Figure 13(a), 13(b), 13(c), and 13(d) that PreDeToR (-$\tau$) is closer to LinUCB and has lower regret than Thomp which also shows that PreDeToR (-$\tau$) is exploiting the latent linear structure of the underlying tasks. In Figure 13(e) we plot the regret of all the baselines with respect to the increasing dimension. Again we see that PreDeToR (-$\tau$) has lower regret than DPT-greedy, AD and Thomp. Observe that with increasing dimension PreDeToR is able to outperform LinUCB. This shows that the PreDeToR (-$\tau$) is able to exploit reward correlation across tasks for varying dimensions.
A.10 Empirical Study: Increasing Attention Heads
@@
Finding 14: PreDeToR (-$\tau$) outperforms DPT-greedy, and AD with increasing attention heads.
@@
Experimental Result: We observe these outcomes in Figure 14. In Figure 14 we show the non-linear bandit setting for horizon $n=20$, $M_{\text{pre}}=160000$, $M_{\text{test}}=200$, $A=20$, $\mathrm{heads}=\{2,4,6,8\}$ and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions atmost once. Observe from Figure 14(a), 14(b), 14(c), and 14(d) that PreDeToR (-$\tau$) has lower regret than AD, Thomp and LinUCB which also shows that PreDeToR (-$\tau$) is exploiting the latent linear structure of the underlying tasks for the non-linear setting. In Figure 14(f) we plot the regret of all the baselines with respect to the increasing attention heads. Again we see that PreDeToR (-$\tau$) regret decreases as we increase the attention heads.
A.11 Empirical Study: Increasing Number of Tasks
@@
Finding 15: PreDeToR (-$\tau$) fails to exploit the underlying latent structure and reward correlation from in-context data when the number of tasks is small.
@@
Experimental Result: We observe these outcomes in Figure 15. In Figure 15 we show the linear bandit setting for horizon $n=20$, $M_{\text{pre}}\in\{5000,10000,50000,100000,150000\}$, $M_{\text{test}}=200$, $A=20$, and $d=40$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$), AD and DPT-greedy suffer more regret than the LinUCB when the number of tasks is small ($M_{\text{train}}\in\{5000,10000\}$ in Figure 15(a), and 15(b). However in Figure 15(c), 15(d), 15(e), and 15(f) we show that PreDeToR has lower regret than Thomp and matches LinUCB. This shows that PreDeToR (-$\tau$) is exploiting the latent linear structure of the underlying tasks for the non-linear setting. Moreover, observe that as $M_{\text{train}}$ increases the PreDeToR has lower cumulative regret than DPT-greedy, AD. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Therefore DPT-greedy does not perform as well as PreDeToR. Finally, note that the result shows that PreDeToR (-$\tau$) is able to exploit the reward correlation across the tasks better as the number of tasks increases.
@@ A.12 Exploration of PreDeToR(-$\tau$) in New Arms Setting
Finding 16: The PreDeToR (-$\tau$) is robust to changes when the number of in-variant actions is large. PreDeToR (-$\tau$) performance drops as shared structure breaks down.
@@
Then in Figure 16(b) we show how the sampling distribution of DPT-greedy, PreDeToR  and PreDeToR-$\tau$ change in the first $10$ and last $10$ rounds for all the tasks where action $17$ is optimal. We plot this graph the same way as discussed in Section 5. From the figure Figure 16(b) we see that PreDeToR(-$\tau$) consistently pulls the action $17$ more than DPT-greedy. It also explores other optimal actions like $\{1,2,3,8,9,15\}$ but discards them quickly in favor of the optimal action $17$ in these tasks.
@@
We now consider the setting where the number of invariant actions is $|\mathcal{A}^{\text{inv}}|=15$. We again show in Figure 18(a) the training distribution of the optimal actions. For each bar, the frequency indicates the number of tasks where the action (shown in the x-axis) is the optimal action. Then in Figure 18(b) we show how the sampling distribution of DPT-greedy, PreDeToR  and PreDeToR-$\tau$ change in the first $10$ and last $10$ rounds for all the tasks where action $17$ is optimal. We plot this graph the same way as discussed in Section 5. From the figure Figure 18(b) we see that none of the algorithms PreDeToR, PreDeToR-$\tau$, DPT-greedy consistently pulls the action $17$ more than other actions. This shows that the common underlying actions across the tasks matter for learning the epxloration.
@@
A.13 Empirical Validation of Theoretical Result
@@
Finding 17: PreDeToR (-$\tau$) excels in predicting the rewards for test tasks when the number of training (source) tasks is large.
(a) Prediction Error for $10^{5}$ tasks
(b) Prediction Error for $2\times 10^{5}$ tasks
@@
A.14 Empirical Study: Offline Performance
@@
Finding 18: PreDeToR (-$\tau$) performs comparably to DPT-greedy and AD in the offline setting.